# Deblurring in the Wild: A Real-World Dataset from Smartphone High-Speed Videos

### v3

@@ 1 Introduction
Image restoration is a widely studied field in computer vision, with research focusing on restoring image quality by removing various forms of noise including blur, haze, noise, and moiréing. Image deblurring refers to the process of restoring the latent sharp image from a blurred image. Motion blur is the most common form of blur in images, and its source is diverse. Fast-moving subjects and/or camera shake can result in motion blur, as well as occlusion in motion boundaries and variation of scene depth.
@@
where II, $B$ and $k$ denote the sharp image, the blurred image and the blur kernel respectively. $\otimes$ represents the convolution operation, and $\eta$ denotes additive noise. The restrictive blur model and assumptions of blur uniformity across image pixels make the estimation of both II and $k$ from $b$ an ill-posed problem, especially when the domain consists of real-life images from various domains.
Later, machine learning based approaches have been developed for the single image deblurring problem. Earlier methods were based on convolutional neural networks (CNNs). They either involved synthesized blur images with a uniform blur kernel for training [11, 12, 13], or were trained to estimate locally linear blur kernels [14]. This is due to the absence of datasets containing real blurry images and corresponding ground truth sharp images as pairs for supervised training. More recent approaches incorporate transformers or state space models. These data-intensive approaches also often struggle to deblur images taken from real-world scenarios and suffer from a loss of generality across blurred images from all domains. For example, some methods based on generative adversarial networks (GANs) only perform well on images containing text or human faces.
The task of building a real-world dataset consisting of blurred and sharp image pairs contains multiple challenges. The primary one is ensuring the geometric and photometric alignment of the blurred image with its corresponding sharp image. The two images have to be taken at the same camera position, which is difficult because camera motion is often essential for introducing blur. There are other necessary criteria for the dataset to be realistic, such as the representation of the most common scenarios for motion blur in the dataset. The presence of noise must also be kept at a minimum level in the ground truth image.
@@
Using iPhone’s 240 fps slow-motion feature, we construct long-exposure blur by averaging frames over a duration (typically 1 second) and use the middle frame as the sharp reference. This mimics physical long-exposure photography while keeping a ground-truth reference. Our dataset comprises 50,000 blur-sharp pairs, significantly larger than existing datasets such as GoPro [15] and RealBlur [16], which contain only a few thousand samples.
@@ 2.1 Image Deblurring Models
The image deblurring problem is addressed in two distinct ways. The first is blind image deblurring, where the goal is to estimate the blur kernel that caused the degradation directly. Classical methods in this domain rely on strong priors or optimization techniques to infer both latent and sharp images, under the assumption of uniform or spatially-varying blur [1, 2, 3, 4, 7, 10, 12, 9, 8, 6, 5]. The second approach involves various deep learning techniques to learn a direct mapping from a blurry image to a sharp one. Deep learning based image deblurring has seen significant advancements in recent years, with the state-of-the-art methods now achieving PSNR scores above 34 dB and SSIM values close to or exceeding 0.97. Leading approaches [17, 18] push the limits of image deblurring by leveraging adaptive reversible decoders and efficient state-space modeling, respectively. A key trend is the growing use of transformer-based architectures [19, 20, 21], which capture long-range dependencies more effectively than traditional CNNs [22, 23, 24]. Another promising direction is the integration of frequency-domain processing, as seen in models like DeepRFT+ [25] and FSNet [26], which exploit the frequency characteristics of blur. Additionally, test-time refinement techniques and lightweight models are actively researched to improve generalization and efficiency in real-world scenarios. Despite these advancements, it is challenging to handle non-uniform blur in real-world dynamic scenes and to generalize across diverse datasets. Emerging works increasingly explore self-supervised learning, Mamba-based state space models, and cross-task unification, pointing to a future where image deblurring becomes more robust and context-aware.
@@ 3.3 Camera Sensor Quality
High-end DSLR cameras and devices like the GoPro HERO11 Black feature larger sensors and hardware-level stabilization, making them less susceptible to motion blur. Smartphones, including the iPhone 15 Pro, have smaller sensors and rely more on computational enhancements. However, this hardware limitation results in more frequent motion blur, which is representative of real-world scenarios. Thus, training models on smartphone-generated blur ensures better generalization to typical user environments.
@@ 3.4 Dataset Diversity and Real-World Relevance
All video sequences and derived image pairs maintain a spatial resolution of 1920×1080 pixels, preserving consistency across samples. The dataset consists of high-quality (blurred, sharp) image pairs directly aligned in space and time. Unlike synthetic datasets that require explicit modeling of the camera response function (CRF) or the application of gamma correction, our dataset avoids such post-processing assumptions. Instead, we extract linear RGB data directly from the iPhone’s video stream, simplifying the data generation pipeline and enhancing physical realism.
@@ 4.1 Benchmarks
Table 1: Performance of SOTA deblurring models on our dataset.
@@
The results shown in Table 1 show that models perform significantly worse than on traditional datasets, demonstrating the difficulty of our benchmark. The average PSNR and SSIM calculated from the input blurry image and the ground truth are 32.38 and 0.777, which sets the baseline performance. None of the papers crossed the baseline PSNR, but FFTFormer exceeded SSIM than the baseline.
Table 2: Comparison of popular image deblurring datasets
@@
Table 2 situates our dataset in the context of existing deblurring benchmarks. Compared to widely used datasets like GoPro and HIDE, our dataset is significantly larger in both the number of image pairs and scene diversity. With over 42,000 pairs across 843 distinct scenes, and a resolution of $1920{\times}1080$, our dataset enables high-fidelity modeling of motion blur. Notably, while it is synthetically generated via frame averaging from high-frame-rate video, the use of consumer-grade smartphone footage provides realistic and diverse blur patterns that reflect real-world camera motion. This combination of scale, diversity, and realism distinguishes our dataset from others, particularly the small-scale, controlled blur datasets like Levin et al. and Köhler et al., and the relatively lower resolution and limited scenes of older benchmarks. Our contribution fills a critical gap in the deblurring literature by offering a modern, challenging dataset that better represents deployment-time scenarios.
@@ 4.2 Visual Examples
To complement our quantitative evaluation, we present a selection of qualitative examples that demonstrate the diversity and complexity of our dataset.
Figure 3: Paired images with sharp and blurred versions.
These samples offer a visual understanding of the types of motion blur captured, as well as the realism of the scenes. By showcasing a range of scenarios, including camera shake, fast-moving objects, and human activity, we aim to illustrate the practical challenges faced by current deblurring models.
@@ 5.1 Limitations
While our dataset represents a significant step forward in realistic motion deblurring benchmarks, it is not without limitations. First, all scenes were recorded under well-lit conditions, primarily during daylight or in well-lit indoor environments, which means low-light scenarios are underrepresented despite being common in real-world photography. Additionally, all benchmark results were obtained using publicly available pretrained models without fine-tuning on our dataset; while this highlights generalization gaps, it may not reflect each model’s full potential performance on our data. The dataset also lacks diversity in capture devices, as all videos were recorded using a single smartphone model. This may limit the generalizability of results to other hardware with different optical or sensor characteristics. Finally, although we report standard quantitative metrics like PSNR and SSIM, we do not include a human perceptual evaluation, which would provide complementary insights into real-world deblurring quality, particularly in complex or ambiguous cases where numerical scores may not correlate with subjective preference.
@@ 5.2 Future Work
Building on the strengths of our dataset, several directions can be pursued to enhance its scope and applicability. One natural extension is to incorporate low-light and nighttime capture conditions, which would expose models to more challenging scenarios involving noise, reduced dynamic range, and sensor non-linearities. Adding data from a broader range of devices including lower-end smartphones, tablets, and action cameras could further diversify the imaging characteristics and improve model robustness across hardware variations. Another promising avenue is to augment the dataset with precise motion metadata, such as gyroscope or optical flow data, which can enable the development of motion-aware or physics-informed deblurring models. Furthermore, collecting user ratings or conducting perceptual studies would help complement traditional metrics like PSNR and SSIM with human-centered evaluations, providing a more comprehensive understanding of visual quality. Finally, leveraging our dataset for training generative or diffusion-based models could open new paths for perceptually grounded and temporally consistent deblurring performance.

### v4

# Deblurring in the Wild: A Real-World Image Deblurring Dataset from Smartphone High-Speed Videos

@@ 1 Introduction
Image restoration is a widely studied field in computer vision, with research focusing on restoring image quality by removing various forms of noise, including blur, haze, noise, and moiréing. Image deblurring refers to the process of restoring the latent sharp image from a blurred image. Motion blur is the most common form of blur in images, and its source is diverse. Fast-moving subjects and/or camera shake can result in motion blur, as well as occlusion in motion boundaries and variation of scene depth.
@@
where $I$, $B$, and $k$ denote the sharp image, the blurred image, and the blur kernel, respectively. $\otimes$ represents the convolution operation, and $\eta$ denotes additive noise. The restrictive blur model and the assumptions of blur uniformity across image pixels make the estimation of both $I$ and $k$ from $b$ an ill-posed problem, especially when the domain consists of real-life images from various domains.
Later, approaches based on machine learning have been developed for the single-image deblurring problem. Earlier methods were based on convolutional neural networks (CNNs). They either involved synthesized blur images with a uniform blur kernel for training [1, 21, 29], or were trained to estimate locally linear blur kernels [24]. This is due to the lack of datasets containing real blurry images and corresponding ground truth sharp images as pairs for supervised training. More recent approaches incorporate transformers or state space models. These data-intensive approaches also often struggle to deblur images taken from real-world scenarios and suffer from a loss of generality across blurred images from all domains. For example, some methods based on generative adversarial networks (GANs) perform well only on images containing text or human faces.
Building a real-world dataset consisting of blurred and sharp image pairs presents multiple challenges. The primary one is ensuring the geometric and photometric alignment of the blurred image with its corresponding sharp image. The two images have to be taken at the same camera position, which is difficult because camera motion is often essential for introducing blur. There are other necessary criteria for the dataset to be realistic, such as the representation of the most common scenarios for motion blur in the dataset. The presence of noise must also be kept at a minimum level in the ground truth image.
@@
Using the iPhone’s 240 fps slow-motion feature, we construct long-exposure blur by averaging frames over a duration (typically 1 second) and use the middle frame as the sharp reference. This mimics physical long-exposure photography while keeping a ground-truth reference. Our dataset comprises 42,000 blur-sharp pairs, significantly larger than existing datasets such as GoPro [18] and RealBlur [20], which contain only a few thousand samples.
@@ 2.1 Image Deblurring Models
The image deblurring problem is addressed in two distinct ways. The first is blind image deblurring, where the goal is to directly estimate the blur kernel that caused the degradation. Classical methods in this domain rely on strong priors or optimization techniques to infer both latent and sharp images, under the assumption of uniform or spatially-varying blur [7, 26, 8, 9, 6, 28, 21, 27, 22, 19, 3]. The second approach involves various deep learning techniques to learn a direct mapping from a blurry image to a sharp one. Such approaches have seen significant advancements in recent years, with the state-of-the-art methods now achieving PSNR scores above 34 dB and SSIM values close to or exceeding 0.97. Leading approaches [16, 12] push the limits of image deblurring by leveraging adaptive reversible decoders and efficient state-space modeling, respectively. A key trend is the growing use of transformer-based architectures [11, 17, 31], which capture long-range dependencies more effectively than traditional CNNs [32, 30, 4]. Another promising direction is the integration of frequency-domain processing, as seen in models like DeepRFT+ [15] and FSNet [5], which exploit the frequency characteristics of blur. Additionally, test-time refinement techniques and lightweight models are actively researched to improve generalization and efficiency in real-world scenarios. Despite these advancements, it is challenging to handle non-uniform blur in real-world dynamic scenes and to generalize across diverse datasets. Emerging works increasingly explore Mamba-based state space models, self-supervised learning, and cross-task unification, pointing to a future where image deblurring becomes more robust and context-aware.
@@ 3 Dataset Construction
All data in our dataset was captured using a single smartphone model, rather than aggregating across multiple devices. We deliberately prioritized benchmark stability and reproducibility over device-agnostic variation. Following precedents set by widely adopted benchmarks such as GoPro and RealBlur, we opted for a fixed hardware setup to ensure controlled blur synthesis, minimize confounding sensor-related variability, and provide a consistent foundation for model evaluation. Modern smartphone cameras from leading manufacturers like Samsung, Pixel, Xiaomi, etc. frequently employ sensors from the Sony IMX family, which share broadly similar imaging characteristics in terms of raw capture and ISP behavior. By contrast, the iPhone’s video system provides unique advantages, particularly its robust hardware-level stabilization and reliable high-frame-rate capture, even up to 240 fps. These features were critical to obtaining temporally aligned, motion-consistent ground truth frames with minimal jitter, which is an essential requirement for high-fidelity blur averaging. Our design focus was to faithfully reproduce real-world degradation patterns stemming from small sensors, low-light conditions, and hand-induced motion blur, rather than variability across devices. The sensor size and frame timing stability were therefore more important to our objectives than manufacturer-specific ISP tuning.
@@
Our blur synthesis pipeline relies on frame averaging to simulate motion blur. While this approach is practical and scalable, it introduces certain simplifications. Specifically, it assumes approximately uniform motion within the averaging window and does not fully reproduce more complex patterns of blur arising from non-linear motion trajectories or occlusions. To approximate realistic worst-case smartphone motion blur, we empirically tested various exposure lengths. Simulating a full 1-second exposure produced excessive saturation and sensor non-linearities, even at minimal ISO levels. We therefore identified  0.5 seconds as the practical upper bound for realistic exposure on modern smartphones. In practice, we averaged 30 consecutive frames at 240 fps, corresponding to an effective blur duration of roughly 1/8 second. This duration offered a principled trade-off: it preserved motion realism while maintaining photometric integrity and sufficient dynamic range. This method assumes simplified motion characteristics, but it provides a physically grounded, reproducible, and scalable synthesis strategy for generating high-fidelity blur across thousands of scenes. Our priority was to establish a large-scale, diverse, and consistent dataset, rather than exhaustively modeling every possible blur phenomenon. Nonetheless, the dataset serves as a robust foundation for future extensions, including non-linear motion modeling, occlusion-aware synthesis, and ablation studies on blur formation.
@@ 3.3 Camera Sensor Quality
High-end DSLR cameras and devices like the GoPro feature larger sensors and hardware-level stabilization, making them less susceptible to motion blur. Smartphones, including the iPhone 15 Pro, have smaller sensors and rely more on computational enhancements. However, this hardware limitation results in more frequent motion blur, which is representative of real-world scenarios. Thus, training models on smartphone-generated blur ensures better generalization to typical user environments.
@@ 3.4 Dataset Diversity and Real-World Relevance
All video sequences and derived image pairs maintain a spatial resolution of 1920×1080 pixels, preserving consistency across samples. The dataset consists of high-quality (blurred, sharp) image pairs directly aligned in space and time. Unlike synthetic datasets that require explicit modeling of the camera response function (CRF) or the application of gamma correction, our dataset avoids such post-processing assumptions. Instead, we extract linear RGB data directly from the iPhone’s video stream, simplifying the data generation pipeline and enhancing physical realism. Our dataset emphasizes well-lit scenarios, a decision driven by the need to maintain high temporal fidelity and spatial alignment when generating blur via frame averaging at 240 fps. In low-light settings, smartphone sensors suffer from elevated noise and reduced dynamic range, which significantly compromise the quality of both the synthesized blur and the sharp reference frame. Prioritizing well-lit scenes ensured high-quality supervision for benchmarking, enabling us to isolate motion-related challenges without conflating them with noise artifacts.
@@ 4.1 Benchmarks
Table 1: Quantitative performance of state-of-the-art deblurring models on our dataset, reported in terms of PSNR and SSIM. The “Blurry vs. GT” row indicates the baseline quality of the input blurry images compared to the ground truth.
@@
Results shown in Table 1 show that all evaluated models perform below this baseline in PSNR, underscoring the difficulty and realism of our benchmark. The average PSNR and SSIM calculated from the input blurry image and the ground truth are 32.38 and 0.777, which sets the baseline performance. None of the papers crossed the baseline PSNR, but FFTFormer and MPRNet exceeded the baseline SSIM.
Table 2: Comparison of widely used image deblurring datasets, highlighting dataset scale, train/test splits, scene diversity, resolution, and data acquisition type. Our dataset provides the largest scale, highest resolution, and broadest scene coverage among existing benchmarks.
@@
Table 2 situates our dataset in the context of existing deblurring benchmarks. Compared to widely used datasets like GoPro and HIDE, our dataset is significantly larger in both the number of image pairs and scene diversity. With over 42,000 pairs across 843 distinct scenes, and a resolution of $1920{\times}1080$, our dataset enables high-fidelity modeling of motion blur. Notably, while it is synthetically generated via frame averaging from high-frame-rate video, the use of consumer-grade smartphone footage provides realistic and diverse blur patterns that reflect real-world camera motion. This combination of scale, diversity, and realism distinguishes our dataset from others, particularly the small-scale, controlled blur datasets like Levin et al. [14] and Köhler et al. [10], and the relatively lower resolution and limited scenes of older benchmarks. Our contribution fills a critical gap in the deblurring literature by offering a modern, challenging dataset that better represents deployment-time scenarios. We report results using PSNR and SSIM for consistency with prior work. However, when others use this dataset to benchmark or evaluate multiple models, human perceptual studies or preference-based evaluation can be incorporated along with standard metrics. This is particularly important for generative approaches, where visual improvements may not be captured by pixel-based measures.
Figure 3: Paired images with sharp and blurred versions.
@@ 4.2 Visual Examples
To complement our quantitative evaluation, we present a selection of qualitative examples that demonstrate the diversity and complexity of our dataset. These samples offer a visual understanding of the types of motion blur captured, as well as the realism of the scenes. By showcasing a range of scenarios, including camera shake, fast-moving objects, and human activity, we aim to illustrate the practical challenges faced by current deblurring models.
@@ 5.1 Limitations
While our dataset represents a significant step forward in realistic motion deblurring benchmarks, it is not without limitations. First, all scenes were recorded under well-lit conditions, primarily during daylight or in well-lit indoor environments, which means low-light scenarios are underrepresented despite being common in real-world photography. Additionally, all benchmark results were obtained using publicly available pretrained models without fine-tuning on our dataset; while this highlights generalization gaps, it may not reflect each model’s full potential performance on our data. The dataset also lacks diversity in capture devices, as all videos were recorded using a single smartphone model. This may limit the generalization of results to other hardware with different optical or sensor characteristics.
@@ 5.2 Future Work
Building on the strengths of our dataset, several directions can be pursued to enhance its scope and applicability. One natural extension is to incorporate low-light and nighttime capture conditions, which would expose models to more challenging scenarios involving noise, reduced dynamic range, and sensor non-linearities. Adding data from a broader range of devices, including lower-end smartphones, tablets, and action cameras, could further diversify the imaging characteristics and improve model robustness across hardware variations. Another promising avenue is to augment the dataset with precise motion metadata, such as gyroscope or optical flow data, which can enable the development of motion-aware or physics-informed deblurring models. Furthermore, collecting user ratings or conducting perceptual studies would help complement traditional metrics like PSNR and SSIM with human-centered evaluations, providing a more comprehensive understanding of visual quality. Finally, leveraging our dataset for training generative or diffusion-based models could open new paths for perceptually grounded and temporally consistent deblurring performance.
Our evaluation relies primarily on PSNR and SSIM, which remain the standard metrics in image restoration benchmarks. While these measures ensure comparability with prior work, they do not fully capture perceptual quality, particularly for generative approaches such as GAN-based models that emphasize realism over pixel-wise accuracy. Metrics like LPIPS or NIQE are better suited to reflect perceptual fidelity, but their inclusion was beyond the scope of this dataset-focused contribution. Our goal in this work was to establish a large-scale, real-world deblurring dataset and to demonstrate its difficulty under widely accepted evaluation protocols. We view the reliance on PSNR/SSIM as a pragmatic choice for benchmarking consistency, while leaving room for future extensions that incorporate perception-oriented metrics for a more comprehensive evaluation of generative methods.